OpenStudy (anonymous):

Simplify square root of 8 divided by fifth root of 8

OpenStudy (anonymous):

i turned them into fractions and got 8 1/2 divided by 8 1/5. but dividing gives me 8 5/2, not one of the answers i can choose from?

OpenStudy (anonymous):

or maybe i'm dividing my fractions wrong...?

OpenStudy (anonymous):

\[\frac{ \sqrt{8} }{ 5\sqrt{8} }\]

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

yeah thats what it looks like

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my options are 9 -1/2 9 -1/4 9 9^2

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

rationalize the denominator